solve
x + 40/x = -13

Answer by checkley77(12844)

x+40/x=-13
(x*x+40)/x=-13
(x^2+40)/x=-13
x^2+40=-13x
x^2+13x+40=0
(x+8)(x+5)=0
x+8=0
x=-8 ans.
x+5=0
x=-5 ans.